OpenStudy (anonymous):

The product of three integers is 90. The second number is twice the first number. The third number is two more than the first number. What are the three numbers?

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.390325.html this shows and explains how the answer was achieved

OpenStudy (anonymous):

let first=x second=2*x third=x+2 x*2x*(x+2)=90 2x^2(x+2)=90 2x^3+4x^2=90 2x^3+4x^2-90=0